# Enhancing Chemistry on Quantum Computers with Fermionic Linear Optical Simulation

### v1

@@ VI-E Nitrogen on the Heron Quantum Processor
These LUCJ circuits were sampled on IBM’s Heron superconducting quantum processor. For this system, we tested subspace dimensions of $\{500^{2},1000^{2},1500^{2},2000^{2}\}$. On average, across all four plots, we again observe reductions in both error and variance for warm-started SQD. As shown in Fig. 9 and Table I, the reductions in error ranged between $32.78\%$ and $46.09\%$ relative to the heat-bath configuration interaction (HCI) energy. The average variance reductions ranged between $33.48\%$ and $82.24\%$, and median variance reductions range between $65.88\%$ and $98.34\%$. In this case, the additional cost of employing ExtraFerm to warm-start SQD was negligible. On average, it required 100 seconds per energy curve, corresponding to an added overhead of at worst $2.03\%$ and at best $42\%.$
@@ VI-F Summary
We observe that ExtraFerm offers substantial performance benefits to SQD through the warm-start approach with very little computational cost. For both molecules, across all subspace dimensions, we see that on average, the warm-started variant of SQD obtained more accurate ground-state energy estimates with lower variance between trials. Most importantly, these results highlight a practical strategy for enhancing the utility of near-term quantum devices for chemistry applications.
[…]
@@ VII Related Works
VII-A Clifford+T Circuits
The Clifford+T gate set, comprising the Clifford group (generated by the phase, Hadamard, and CNOT gates) and the T gate, is among the most extensively studied in quantum computing. Stabilizer circuits, which contain only Clifford gates, are classically simulable in polynomial time by the Gottesman-Knill theorem [gottesman1998heisenberg]. Adding T gates enables universal quantum computation [bravyi2005universal]. States produced by a small number of T gates have low stabilizer rank and can be expressed as a concise linear combination of stabilizer states [bravyi2019simulation], making them amenable to classical simulation. Ref. [bravyi2016improved] leverages this property to propose a Monte Carlo probability estimation routine for Clifford-dominated circuits. This algorithm is similar in spirit to Reardon-Smith’s [reardon2024improved], which ExtraFerm builds upon.
VII-B Tensor Networks
Tensor networks [white1992density, orus2014practical] are a framework for representing quantum many-body systems and are a popular choice for simulating quantum circuits [markov2008simulating]. Tensor networks factorize quantum states into smaller tensors, which works particularly well in instances for which there is limited entanglement. However, this method may scale poorly for some quantum circuits with many nonlocal gates.
@@ VII-C Classical Simulation to Boost Quantum Fidelity
Recently, several works have used classical simulation to mitigate errors on quantum computers. These methods typically simulate a subset of the quantum system or an efficiently simulable related system. For instance, Clifford Data Regression [czarnik2021error] fits a linear model using Clifford-dominated circuits to map noisy observables to their noiseless counterparts, achieving error reductions between one and two orders of magnitude. Liu et al. [liu2022classical] offload the computation of an auxiliary error-mitigation circuit to a classical processor, further reducing noise effects. Finally, CAFQA [ravi2022cafqa] accelerates convergence of variational quantum algorithms by searching over Clifford circuits to identify high-quality initial parameters, yielding substantial accuracy gains.
VIII Discussion
As the utility of near-term quantum computers increases, researchers need tools to assess the quality of data collected on noisy quantum hardware. We have introduced ExtraFerm, a simulator that can be used to estimate Born-rule probabilities of bitstrings with respect to particle number-conserving extended matchgate circuits. We have seen that, in addition to being a useful tool for simulating chemistry circuits, ExtraFerm can be integrated with SQD to improve the quality of molecular ground-state energy estimates through a warm-start technique. We propose two directions for future work in which ExtraFerm could further advance the practical application of quantum chemistry on current quantum hardware.
First, ExtraFerm could enable sophisticated post-processing strategies for noisy samples from quantum hardware. While we found that the simple approach of selecting high-probability bitstrings during early iterations of SQD improved energy estimates, there are a variety of other techniques that one could try. The ability to compute Born-rule probabilities of existing data could provide a way to re-weight the observed bitstrings and re-sample from the pool of collected data. Discrepancies between the empirical and simulated distributions could also be used to infer or correct aspects of the underlying noise model of the device. In this way, ExtraFerm may support powerful error-mitigation strategies for chemistry circuits on near-term quantum devices.
Second, rather than using ExtraFerm to post-process measurements obtained from noisy quantum devices, one could employ it to optimize the parameters of a quantum circuit before it is executed on hardware. For instance, in a quantum chemistry setting, suppose one has an approximate ansatz that prepares a target electronic state. Given access to estimates of amplitudes in the target state, ExtraFerm could be used to minimize the discrepancy between the target probabilities of selected bitstrings (corresponding to important configurations) and the probabilities produced by the circuit. Gradient-free methods such as simultaneous perturbation stochastic approximation (SPSA) [spall2002implementation] could be used to facilitate optimization.
We hope that researchers working with matchgate-based chemistry circuits will find ExtraFerm useful and discover new ways of incorporating it into their workflows. Furthermore, we hope that the spirit of our integration of ExtraFerm with SQD will encourage the development of more hybrid approaches that use classical computation to extend the capabilities of quantum chemistry on near-term quantum hardware.
@@ Acknowledgments
The calculations in Section VI were performed on the Midway3 Cluster, maintained by the University of Chicago Research Computing Center. We obtained the nitrogen dataset of Ref. [robledo2025sqd] from https://github.com/jrm874/sqd_data_repository.
-A Proof of Exact Calculation
$p_{\text{true}}$ $=\left|\left\langle b\right|V_{k+1}\prod_{j=1}^{k}C(\theta_{j})V_{j}\left|a\right\rangle\right|^{2}$ (20)
$=\left|\left\langle b\right|V_{k+1}\prod_{j=1}^{k}\left(\cos\left(\frac{\theta_{j}}{4}\right)D_{0}(\theta_{j})+i\sin\left(\frac{\theta_{j}}{4}\right)D_{1}(\theta_{j})\right)V_{j}\left|a\right\rangle\right|^{2}$ (21)
$=\left|\left\langle b\right|\sum_{x\in\{0,1\}^{k}}V_{k+1}\prod_{j=1}^{k}\left(\cos\left(\frac{\theta_{j}}{4}\right)D_{0}(\theta_{j})\right)^{x_{j}-1}\cdot\left(i\sin\left(\frac{\theta_{j}}{4}\right)D_{1}(\theta_{j})\right)^{x_{j}}V_{j}\left|a\right\rangle\right|^{2}$ (22)
$=\left|\sum_{x\in\{0,1\}^{k}}\left(\prod_{j=1}^{k}\cos\left(\frac{\theta_{j}}{4}\right)^{x_{j}-1}\cdot\left(i\sin\left(\frac{\theta_{j}}{4}\right)\right)^{x_{j}}\right)\left\langle b\right|V(x)\left|a\right\rangle\right|^{2}$ (23)

### v2

@@ VI-E Nitrogen on the Heron Quantum Processor
These LUCJ circuits were sampled on IBM’s Heron superconducting quantum processor. For this system, we tested subspace dimensions of $\{500^{2},1000^{2},1500^{2},2000^{2}\}$. On average, across all four plots, we again observe reductions in both error and variance for warm-started SQD. As shown in Fig. 10 and Table I, the reductions in error ranged between $32.78\%$ and $46.09\%$ relative to the heat-bath configuration interaction (HCI) energy. The average variance reductions ranged between $33.48\%$ and $82.24\%$, and median variance reductions range between $65.88\%$ and $98.34\%$. The additional cost of employing ExtraFerm to warm-start SQD was negligible. On average, it required 100 seconds per energy curve, corresponding to an added overhead of at worst $2.03\%$ and at best $42\%.$
[…]
@@ VI-F Summary
We observe that ExtraFerm offers substantial performance benefits to SQD through the warm-start approach with very little computational overhead. For both systems that we study, across all subspace dimensions, we see that on average, the warm-started variant of SQD obtained more accurate ground-state energy estimates with lower variance between trials. Most importantly, these results highlight a practical strategy for enhancing the utility of near-term quantum devices for chemistry applications.
@@ VII Related Works
VII-A Simulation of Clifford+T Circuits
The Clifford+T gate set, comprising the Clifford group (generated by the phase, Hadamard, and CNOT gates) and the T gate, is among the most extensively studied in quantum computing. Stabilizer circuits, which contain only Clifford gates, are classically simulable in polynomial time by the Gottesman-Knill theorem [gottesman1998heisenberg]. Adding T gates enables universal quantum computation [bravyi2005universal]. States produced by a small number of T gates have low stabilizer rank and can be expressed as a concise linear combination of stabilizer states [bravyi2019simulation], making them amenable to classical simulation. Bravyi and Gosset [bravyi2016improved] leverage this property to introduce a Monte Carlo probability estimation routine for Clifford-dominated circuits. This algorithm is similar in spirit to Reardon-Smith’s [reardon2024improved] (which ExtraFerm builds upon), though it is tailored to a different class of circuits.
VII-B Other Simulation Frameworks
Tensor networks [white1992density, orus2014practical] are a framework for representing quantum many-body systems and are a popular choice for simulating quantum circuits [markov2008simulating]. They express quantum states as a collection of complex-valued multidimensional arrays. However, tensor network methods may scale poorly for systems with high entanglement [berezutskii2025tensor]. Recent work has also simulated the LUCJ ansatz using tensor networks [rudolph2025simulating].
Alternatively, Pauli propagation [rudolph2025paulipropagationcomputationalframework] methods approximate the evolution of a quantum operator by truncating the Pauli path integral, enabling fast, rough estimates of expectation values even on arbitrary circuit topologies. Initially shown to be effective in noisy and random settings, the approach has since been applied more broadly and appears promising for variational use cases. It is also well suited for hybrid quantum–classical workflows. Pauli propagation simulation costs scale exponentially with non-Clifford gates and accuracy requirements.
@@ VII-C Classical Simulation to Boost Quantum Fidelity
Recently, several works have used classical simulation to mitigate errors on quantum computers. These methods typically simulate a subset of the quantum system or an efficiently simulable related system. For instance, Clifford Data Regression [czarnik2021error] fits a linear model using Clifford-dominated circuits to map noisy observables to their noiseless counterparts, achieving error reductions between one and two orders of magnitude. Ref. [liu2022classical] offloads the computation of an auxiliary error-mitigation circuit to a classical processor, further reducing noise effects. Finally, CAFQA [ravi2022cafqa] accelerates convergence of variational quantum algorithms by searching over Clifford circuits to identify high-quality initial parameters, yielding substantial accuracy gains.
VIII Conclusion and Discussion
We have introduced ExtraFerm, a quantum circuit simulator tailored to particle number-conserving matchgates and controlled-phase gates. We have seen that ExtraFerm is capable of significant advantages in latency performance and exponential advantages in memory performance memory over tensor network and state vector methods in certain regimes. Most notably, we propose a method for integrating ExtraFerm with SQD to improve the quality of molecular ground-state energy estimates using quantum computers with almost no computational overhead. We anticipate two future directions in which ExtraFerm may further aid quantum-classical workflows for chemistry.
First, ExtraFerm’s ability to compute approximate Born-rule probabilities could be used to enable sophisticated post-processing strategies for noisy quantum data including importance sampling [tokdar2010importance]. Second, one could use ExtraFerm to optimize the parameters of a quantum circuit by minimizing the discrepancies between the ideal Born-rule probabilities of a set of measurements and those given by the circuit. Gradient-free methods such as simultaneous perturbation stochastic approximation (SPSA) [spall2002implementation] could be used to facilitate this optimization.
We hope that researchers working with matchgate-based chemistry circuits will find ExtraFerm useful and discover new ways of incorporating it into their workflows. Furthermore, we expect that the spirit of our integration of ExtraFerm with SQD will encourage the development of more hybrid approaches that use classical computation to extend the capabilities of quantum chemistry on near-term hardware.
@@ Acknowledgments
The calculations in Section VI were performed on the Midway3 Cluster, maintained by the University of Chicago Research Computing Center. We obtained the nitrogen dataset used in Ref. [robledo2025sqd] from https://github.com/jrm874/sqd_data_repository.